### v1

# The exact travelling wave solutions of a KPP equation

Abstract
We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed. The solutions are written in terms of elementary functions.
I Introduction
Fisher fisher first introduced a nonlinear evolution equation to investigate the wave propagation of an advantageous gene in a population. His equation also describes the logistic growth–diffusion process and has the form
$u_{t}-\nu u_{xx}=ku\left(1-\frac{u}{\kappa}\right)$ (1)
where $\kappa(>0)$ is the carrying capacity of the environment. The term $f(u)=ku(1-u/\kappa)$ represents a nonlinear growth rate which is proportional to $u$ for small $u$, but decreases as $u$ increases, and vanishes when $u=\kappa$. Equation (1) arises in many physical, biological, and chemical problems involving diffusion and nonlinear growth. It is convenient to introduce the nondimensional quantities $x^{*},t^{*},u^{*}$, so after dropping the asterisks, equation (1) takes the nondimensional form
$u_{t}-u_{xx}=u\left(1-u\right).$ (2)
Fisher equation admits traveling wave solutions
$u(x,t)=u(\xi),\hskip 14.22636pt\xi=x-ct,$ (3)
where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=u\left(u-1\right).$ (4)
The aim of the present paper is to show that the generalization of Eq. (4), to be presented below, can be integrated in elementary functions for the specific values of the speed $c$.
In the same year (1937) as Fisher, Kolmogorov, Petrovsky and Piskunov (KPP) kolmogorov introduced the more general reaction-diffusion equation, which in the nondimensional form can be presented as
$u_{t}-u_{xx}=F(u),$ (5)
where $F$ is a sufficiently smooth function with the properties that $F(0)=F(1)=0$, $F^{\prime}(0)>0$ and $F^{\prime}(1)<0$.
II Travelling waves
We consider the equation
$u_{t}-u_{xx}=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (6)
where $n$ and $a$ are some positive constants. Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (6) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath . For the solutions of (6) in the form (3), $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right)$ (7)
with the boundary conditions (11). Equation (7) has a simple mechanical interpretation. It describes motion of the Newtonian particle with the mass 1 ($u$ being the coordinate of the particle and $\xi$ the time) in the potential well $U(u)$ defined by the equation
$\frac{dU}{du}=-u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (8)
The motion is with friction, and the coefficient of friction is equal to $c$. The point $u=0$ is the point of stable equilibrium, and the point $u=1$ unstable.
Equation (7) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
$p=\frac{du}{d\xi}$ (9)
as the new dependent variable. In the new variables (7) takes the form of Abel equation of the second kind polyanin .
$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (10)
The boundary conditions for (7) were
$\lim_{\xi\to-\infty}u(\xi)$ $=1,$ (11a)
$\lim_{\xi\to+\infty}u(\xi)$ $=0.$ (11b)
The boundary conditions for (10) are
$p(1)$ $=1,$ (12a)
$p(0)$ $=0.$ (12b)
One can easily check up that for
$c=\frac{(n+1)a+1}{\sqrt{n+1}},$ (13)
the solution of (10) satisfying the boundary conditions (12) is
$p=\frac{u\left(u^{n}-1\right)}{\sqrt{n+1}}.$ (14)
Substituting $p(u)$ into (9) and integrating we obtain the solution of (7) as
$u(\xi)=\frac{1}{\left[\exp\left(n\xi/\sqrt{n+1}\right)+1\right]^{1/n}}.$ (15)
For the Fisher equation we obtain the well known result ablo
$u(\xi)=\frac{1}{\left[\exp\left(\xi/\sqrt{6}\right)+1\right]^{2}},$ (16)
and the condition (13) takes the form $c=5/\sqrt{6}$.
III Perturbation series
In the previous Section we used the method of integration based on Eq. (10) to find the exact integral of Eq. (7) for the specific value of $c$. Now we want to consider the limiting case $c\gg 1$ and show how the former equation can be used to construct the perturbation series for the solution of the latter.
Expanding the solution of (10) with respect to the parameter $1/c$
$p(u)=\sum_{k=1}\frac{1}{c^{k}}p^{(k)}(u),$ (17)
we obtain
$p^{(1)}$ $=u\left(u^{n}-1\right)\left(u^{n}+a\right),$ (18a)
$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u}.$ (18b)
To make our life simpler, further on consider the particular case of (10) corresponding to the Fisher equation:
$pp_{u}+cp=u(u-1).$ (19)
In this case the first two terms of the expansion (17) are
$p^{(1)}$ $=u(u-1),$ (20a)
$p^{(2)}$ $=u(u-1)(1-2u).$ (20b)
Substituting $p=p^{(1)}+p^{(2)}/c$ into (9) we obtain after a bit of algebra debnath
$u(\xi)$ $=$ $\frac{1}{\exp\left(\xi/c\right)+1}$ (21)
$+$ $\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$
IV Discussion
Let us return to (10). We have the first order ODE with two boundary conditions (12). The first impression is that the problem is overdetermined, and the solution obtained in Section II exists due to the finetuning of the l.h.s. of the equation (by choosing $c$ given by Eq. (13)). (It is irrelevant for the time being that the solution both of (10) and (7) turn out to be the elementary functions.) And the solution of (10) will not exist for any other value of $c$. However, the last statement immediately comes into contradiction with the results of the previous Section. So what about our first impression?
Notice that the boundary conditions (12) for Eq. (10) are exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$. So the boundary conditions should be taken with care. In what follows, we’ll understand that all boundary conditions are equal, but some are more equal than the others.
Let us return to Eq. (7) and consider the solution in the vicinity of $u=0$ and $u=1$. In the first case, the equation can be presented as
$u_{\xi}$ $=p,$ (22a)
$p_{\xi}$ $=-cp-au.$ (22b)
In the second case, the equation can be presented as
$v_{\xi}$ $=-p,$ (23a)
$p_{\xi}$ $=cp-(a+1)nv,$ (23b)
where $v=1-u$.
Let us start from analysing Eq. (23). The characteristic equation for the system is
$s^{2}-cs-n(a+1)=0,$ (24)
the roots being
$s_{1,2}$ $=\frac{c\pm\sqrt{c^{2}+4n(a+1)}}{2}.$ (25)
(Further on $s_{1}$ will be the positive root, and $s_{2}$ the negative.) Hence the system can be reduced to the canonical form
$X_{\xi}$ $=s_{1}X,$ (26a)
$Y_{\xi}$ $=s_{2}Y,$ (26b)
where
$X$ $=p+s_{2}v,$ (27a)
$Y$ $=p+s_{1}v.$ (27b)
Because $u(\xi)$ should satisfy the boundary condition (11a), we obtain
$Y=0.$ (28)
Returning to Eq. (10) we realize that the boundary condition (12a) means that in the vicinity of $u=1$
$p(u)=s_{1}(u-1).$ (29)
Now let us come to (22). The characteristic equation for the system is
$s^{2}+cs+a=0.$ (30)
The roots of the characteristic equations are
$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4a}}{2}.$ (31)
We have to consider separately the case of the real roots, the case of the complex roots and the case of the double root.
Let us start from the case of the real roots ($c>2\sqrt{a}$). In this case we can reduce the system (22) to
$\frac{dX}{s_{3}X}=\frac{dY}{s_{4}Y},$ (32)
where
$X$ $=p+s_{4}v,$ (33a)
$Y$ $=p+s_{3}v.$ (33b)
The general integral is given by the equation
$Y=CX^{\frac{s_{2}}{s_{1}}}.$ (34)
Because $s_{3}$ and $s_{4}$ have the same sign, $X$ approaches zero with $Y$, and all the integral curves pass through the origin, which is a node. (Note that $c$ given by (13) obviously fulfills the inequality $c>2\sqrt{a}$ for any $n>0$ and $a>0$.)
For $c<2\sqrt{a}$ the characteristic equation has two conjugate imaginary roots $\alpha+\beta i,\alpha-\beta i$, $\beta\neq 0$) we can reduce the system (22) to the form goursat
$\frac{dX}{\alpha X-\beta Y}=\frac{dY}{\beta X+\alpha Y}$ (35)
from which we derive
$\frac{XdX+YdY}{\alpha\left(X^{2}+Y^{2}\right)}=\frac{XdY-YdX}{\beta\left(X^{2}+Y^{2}\right)}$ (36)
The general integral of the equation (122) is therefore represented by the equation
$\sqrt{X^{2}+Y^{2}}=Ce^{\frac{\alpha}{\beta}\tan^{-1}\frac{Y}{X}}.$ (37)
All these curves have the form of spirals which approach the origin as an asymptotic point. The origin is said to he a focus.
If the characteristic equation has a double root $s=c/2$, the system (22) reduces to the form
$\frac{dX}{X}=\frac{dY}{X+Y},$ (38)
and the general integral is $Y=CX+X\ln X$. In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$. When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus.
Now we can resolve the (seeming) contradiction mentioned in the beginning of the present Section. Actually we have to solve ODE with a single boundary condition (12a) (presented as (29)). The second boundary condition (12b) is satisfied automatically

### v2

@@
 $u_{t}-u_{xx}=u\left(1-u\right).$ (2)
-Fisher equation admits traveling wave solutions
-$u(x,t)=u(\xi),\hskip 14.22636pt\xi=x-ct,$ (3)
-where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
-$u_{\xi\xi}+cu_{\xi}=u\left(u-1\right).$ (4)
-The aim of the present paper is to show that the generalization of Eq. (4), to be presented below, can be integrated in elementary functions for the specific values of the speed $c$.
 In the same year (1937) as Fisher, Kolmogorov, Petrovsky and Piskunov (KPP) kolmogorov introduced the more general reaction-diffusion equation, which in the nondimensional form can be presented as
@@
 II Travelling waves
-We consider the equation
-$u_{t}-u_{xx}=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (6)
-where $n$ and $a$ are some positive constants.
-Note that for $n=1/2$ and $a=1$ we recover Fisher equation.
-Equation (6) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
-For the solutions of (6) in the form (3), $u(\xi)$ satisfies the equation
-$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right)$ (7)
-with the boundary conditions (11).
+Equation (3) admits traveling wave solutions
+$u(x,t)=u(\xi),\hskip 14.22636pt\xi=x-ct,$ (4)
+where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
+$u_{\xi\xi}+cu_{\xi}=F(u)$ (5)
+with the boundary conditions
+$\lim_{\xi\to-\infty}u(\xi)$ $=1,$ (6a)
+$\lim_{\xi\to+\infty}u(\xi)$ $=0.$ (6b)
 Equation (5) has a simple mechanical interpretation.
 It describes motion of the Newtonian particle with the mass 1 ($u$ being the coordinate of the particle and $\xi$ the time) in the potential well $U(u)$ defined by the equation
-$\frac{dU}{du}=-u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (8)
+$\frac{dU}{du}=-F(u).$ (7)
 The motion is with friction, and the coefficient of friction is equal to $c$. The point $u=0$ is the point of stable equilibrium, and the point $u=1$ unstable.
+III The ODE which doesn’t contain explicitly the independent variable
 Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
@@
 as the new dependent variable. In the new variables (5) takes the form of Abel equation of the second kind polyanin .
-$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (10)
-The boundary conditions for (7) were
-$\lim_{\xi\to-\infty}u(\xi)$ $=1,$ (11a)
-$\lim_{\xi\to+\infty}u(\xi)$ $=0.$ (11b)
-The boundary conditions for (10) are
+$pp_{u}+cp=F(u).$ (9)
+The boundary conditions for (5) were (6). The boundary conditions for (9) are (however, see the next Section)
 $p(1)$ $=1,$ (10a)
 $p(0)$ $=0.$ (10b)
-One can easily check up that for
-$c=\frac{(n+1)a+1}{\sqrt{n+1}},$ (13)
-the solution of (10) satisfying the boundary conditions (12) is
-$p=\frac{u\left(u^{n}-1\right)}{\sqrt{n+1}}.$ (14)
-Substituting $p(u)$ into (9) and integrating we obtain the solution of (7) as
-$u(\xi)=\frac{1}{\left[\exp\left(n\xi/\sqrt{n+1}\right)+1\right]^{1/n}}.$ (15)
+Now let us be more specific and consider $F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right)$, where $n>0$ and $a\geq 0$ are some constants, so Eqs. (5) and (9) respectively take the form
+$u_{\xi\xi}+cu_{\xi}$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (11a)
+$pp_{u}+cp$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (11b)
+Note that for $n=1/2$ and $a=1$ we recover Fisher equation.
+Equation (11a) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
+One can easily check up that there exists the solution of (11b) satisfying the boundary conditions (10)
+$p=Pu\left(u^{n}-1\right),$ (12)
+where
+$P=\frac{1}{\sqrt{n+1}},$ (13)
+provided that the speed of the wave is
+$c=P+\frac{a}{P}.$ (14)
+Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (11a) as
+$u(\xi)=\frac{1}{\left[\exp(nP\xi)+1\right]^{1/n}}.$ (15)
 For the Fisher equation we obtain the well known result ablo
@@
 and the condition (14) takes the form $c=5/\sqrt{6}$.
-III Perturbation series
-In the previous Section we used the method of integration based on Eq.
-(10) to find the exact integral of Eq.
-(7) for the specific value of $c$.
-Now we want to consider the limiting case $c\gg 1$ and show how the former equation can be used to construct the perturbation series for the solution of the latter.
-Expanding the solution of (10) with respect to the parameter $1/c$
-$p(u)=\sum_{k=1}\frac{1}{c^{k}}p^{(k)}(u),$ (17)
-we obtain
-$p^{(1)}$ $=u\left(u^{n}-1\right)\left(u^{n}+a\right),$ (18a)
-$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u}.$ (18b)
-To make our life simpler, further on consider the particular case of (10) corresponding to the Fisher equation:
-$pp_{u}+cp=u(u-1).$ (19)
-In this case the first two terms of the expansion (17) are
-$p^{(1)}$ $=u(u-1),$ (20a)
-$p^{(2)}$ $=u(u-1)(1-2u).$ (20b)
-Substituting $p=p^{(1)}+p^{(2)}/c$ into (9) we obtain after a bit of algebra debnath
-$u(\xi)$ $=$ $\frac{1}{\exp\left(\xi/c\right)+1}$ (21)
-$+$ $\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$
-IV Discussion
-Let us return to (10).
-We have the first order ODE with two boundary conditions (12).
-The first impression is that the problem is overdetermined, and the solution obtained in Section II exists due to the finetuning of the l.h.s.
-of the equation (by choosing $c$ given by Eq.
-(13)).
-(It is irrelevant for the time being that the solution both of (10) and (7) turn out to be the elementary functions.) And the solution of (10) will not exist for any other value of $c$.
-However, the last statement immediately comes into contradiction with the results of the previous Section.
-So what about our first impression?
-Notice that the boundary conditions (12) for Eq.
-(10) are exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
-So the boundary conditions should be taken with care.
-In what follows, we’ll understand that all boundary conditions are equal, but some are more equal than the others.
-Let us return to Eq. (7) and consider the solution in the vicinity of $u=0$ and $u=1$. In the first case, the equation can be presented as
+In Appendix A we’ll see how the results (12) - (15) can be generalized in the case when advection is taken into account in addition to diffusion.
+IV The correct boundary conditions for Eq. (9)
+In the previous Section we postulated two boundary conditions (10) for (9).
+However, the equation is of the first order.
+The first impression is that the problem is overdetermined, and the solution obtained there exists due to the finetuning of the l.h.s.
+of the equation.
+So one may think that we should finetune the speed $c$ for the solution of (11b) to exist, and the solution will not exist for a general value of $c$.
+In fact, didn’t we finetune $c$ in the previous Section?
+However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (11b) for arbitrary $c\gg 1$ is presented.
+It looks like we have a problem.
+We have another problem with the boundary conditions (10) - they are exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
+In this Section we’ll solve both problems.
+We’ll show that one of the boundary conditions (10) should be specified, and the other can be discarded.
+Consider the solution of Eq. (5) in the vicinity of $u=0$ and $u=1$. In the first case, the equation can be presented as
 $u_{\xi}$ $=p,$ (17a)
-$p_{\xi}$ $=-cp-au.$ (22b)
+$p_{\xi}$ $=-cp-F^{\prime}(0)u.$ (17b)
 In the second case, the equation can be presented as
 $v_{\xi}$ $=-p,$ (18a)
-$p_{\xi}$ $=cp-(a+1)nv,$ (23b)
+$p_{\xi}$ $=cp-|F^{\prime}(1)|v,$ (18b)
 where $v=1-u$.
-Let us start from analysing Eq. (23). The characteristic equation for the system is
-$s^{2}-cs-n(a+1)=0,$ (24)
-the roots being
-$s_{1,2}$ $=\frac{c\pm\sqrt{c^{2}+4n(a+1)}}{2}.$ (25)
-(Further on $s_{1}$ will be the positive root, and $s_{2}$ the negative.) Hence the system can be reduced to the canonical form
+Let us start from analysing (18). The system can be reduced to the canonical form
 $X_{\xi}$ $=s_{1}X,$ (19a)
@@
 $X$ $=p+s_{2}v,$ (20a)
-$Y$ $=p+s_{1}v.$ (27b)
+$Y$ $=p+s_{1}v,$ (20b)
+and $s_{1,2}$ are the roots of the characteristic equation
+$s^{2}-cs-|F^{\prime}(1)|=0,$ (21)
+that is
+$s_{1,2}$ $=\frac{c\pm\sqrt{c^{2}+4|F^{\prime}(1)|}}{2}.$ (22)
+(Further on $s_{1}$ will be the positive root, and $s_{2}$ the negative.)
 Because $u(\xi)$ should satisfy the boundary condition (6a), we obtain
 $Y=0.$ (23)
-Returning to Eq. (10) we realize that the boundary condition (12a) means that in the vicinity of $u=1$
-$p(u)=s_{1}(u-1).$ (29)
+Returning to Eq. (9) we realize that the boundary condition (10a) should be specified as follows: In the vicinity of $u=1$
+$p=s_{1}(u-1).$ (24)
+For practical purposes we can sometimes continue to use this boundary condition in the previous form (10a), like we deed in Section III.
+The differential equation itself produces the correct condition (24).
+However, if we want, for example, to solve (9) numerically, the condition (24) should be used from the beginning.
 Now let us come to (17). The characteristic equation for the system is
-$s^{2}+cs+a=0.$ (30)
+$s^{2}+cs+F^{\prime}(0)=0.$ (25)
 The roots of the characteristic equations are
-$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4a}}{2}.$ (31)
+$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4F^{\prime}(0)}}{2}.$ (26)
 We have to consider separately the case of the real roots, the case of the complex roots and the case of the double root.
-Let us start from the case of the real roots ($c>2\sqrt{a}$). In this case we can reduce the system (22) to
+Let us start from the case of the real roots ($c^{2}>4F^{\prime}(0)$). In this case we can reduce the system (17) to
 $\frac{dX}{s_{3}X}=\frac{dY}{s_{4}Y},$ (27)
@@
 The general integral is given by the equation
-$Y=CX^{\frac{s_{2}}{s_{1}}}.$ (34)
+$Y=CX^{\frac{s_{4}}{s_{3}}}.$ (29)
 Because $s_{3}$ and $s_{4}$ have the same sign, $X$ approaches zero with $Y$, and all the integral curves pass through the origin, which is a node.
-(Note that $c$ given by (13) obviously fulfills the inequality $c>2\sqrt{a}$ for any $n>0$ and $a>0$.)
-For $c<2\sqrt{a}$ the characteristic equation has two conjugate imaginary roots $\alpha+\beta i,\alpha-\beta i$, $\beta\neq 0$) we can reduce the system (22) to the form goursat
+(Note that $c$ given by (14) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
+For $c^{2}<4F^{\prime}(0)$ the characteristic equation has two conjugate imaginary roots $\alpha+\beta i,\alpha-\beta i$, $\beta\neq 0$) we can reduce the system (17) to the form goursat
 $\frac{dX}{\alpha X-\beta Y}=\frac{dY}{\beta X+\alpha Y}$ (30)
@@
 All these curves have the form of spirals which approach the origin as an asymptotic point. The origin is said to he a focus.
-If the characteristic equation has a double root $s=c/2$, the system (22) reduces to the form
+If the characteristic equation has a double root $s=c/2$, which occurs for $c^{2}=4F^{\prime}(0)$, the system (17) reduces to the form
 $\frac{dX}{X}=\frac{dY}{X+Y},$ (33)
 and the general integral is $Y=CX+X\ln X$.
-In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$.
+In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e^{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$.
 When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus.
-Now we can resolve the (seeming) contradiction mentioned in the beginning of the present Section.
-Actually we have to solve ODE with a single boundary condition (12a) (presented as (29)).
-The second boundary condition (12b) is satisfied automatically
+Note that if for physical reasons $u$ should be strictly nonnegative, the focus solutions (and hence the value of speed $c$ for which $c^{2}\leq 4F(0)$) are unacceptable.
+Now we understand that the second boundary condition (10b) can be discarded.
+It is satisfied automatically.
+And returning to the previous Section: we did finetune the speed $c$ for the solution of (11b) to be expressed in terms of elementary functions, and not for the solution to exist.
+Appendix A Generalized Burgers–Huxley equation
+The results obtained in Section II for the generalized KPP equation (11a) can be easily generalised to the generalized Burgers–Huxley equation cherniha ; kushner
+$u_{t}-\lambda u^{n}u_{x}-u_{xx}=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (34)
+where $\lambda$ is a constant.
+(The Burgers–Huxley equation is known in various fields of applied mathematics.
+For example, it describes transport processes in systems when diffusion and convection are equally important and nonlinear reaction–diffusion processes too.)
+For the travelling wave solution of (34) we obtain instead of (11b) the equation
+$pp_{u}+\lambda u^{n}p+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (35)
+One can easily see that the solution (12) (and hence Eq. (15)) is valid in the present case also, same as the formula (14). The only difference is that Eq. (13) in this case takes the form
+$P=\frac{\sqrt{\lambda^{2}+4(n+1)}-\lambda}{2(n+1)}.$ (36)
+If we put $a=0$ in the formula (14), Eqs. (36) and (15) supply the solution of the generalized Huxley equation clarkson
+$u_{t}-\lambda u^{n}u_{x}-u_{xx}=u^{n+1}\left(1-u^{n}\right).$ (37)
+Appendix B Perturbation series
+Let us return to Eq.
+(9).
+In this Appendix we’ll present the expansion of the solution of the equation (and hence of the solution of Eq.
+(5)) with respect to the powers of $1/c$ (meaningful if $c\gg 1$).
+Such expansion has the form
+$p(u)=\sum_{k=1}\frac{1}{c^{k}}p^{(k)}(u),$ (38)
+where
+$p^{(1)}$ $=F(u),$ (39a)
+$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u}.$ (39b)
+Let us restrict ourselves by the next to leading order approximation with respect to the parameter $1/c$. Substituting thus truncated $p(u)$ into (8) we obtain
+$\frac{du}{d\xi}=\frac{1}{c}F(u)\left(1-\frac{1}{c}\frac{dF(u)}{du}\right)$ (40)
+or, equivalently (with the accepted precision),
+$\frac{d\xi}{c}=\frac{du}{F(u)}+\frac{1}{c}\frac{dF(u)}{F(u)}.$ (41)
+Integrating (41) we get
+$\frac{\xi}{c}=\int\frac{du}{F(u)}+\frac{1}{c}\ln|F(u)|.$ (42)
+To make our life simpler, further on consider only the particular case of $F(u)$ corresponding to the Fisher equation:
+$F(u)=u(u-1).$ (43)
+Performing integration in (42) we obtain
+$\frac{\xi}{c}=\ln\left[\frac{1-u}{u}\right]+\frac{1}{c}\ln\left[u(1-u)\right].$ (44)
+In the leading order with respect to $1/c$ the solution of (44) is
+$u(\xi)=$ $\frac{1}{\exp\left(\xi/c\right)+1},$ (45)
+hence with the accepted precision we can rewrite (44) as
+$\frac{\xi}{c}-\frac{1}{c}\ln\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}=\ln\left[\frac{1-u}{u}\right].$ (46)
+Exponentiating, solving with respect to $u$ and again expanding with respect to $1/c$ we obtain after a bit of algebra
+$u(\xi)=\frac{1}{\exp\left(\xi/c\right)+1}$
+$+\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$ (47)

### v3

@@
 III The ODE which doesn’t contain explicitly the independent variable
-Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
+Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea benguria ; kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
 $p=\frac{du}{d\xi}$ (8)
@@
 The boundary conditions for (5) were (6). The boundary conditions for (9) are (however, see the next Section)
-$p(1)$ $=1,$ (10a)
+$p(1)$ $=0,$ (10a)
 $p(0)$ $=0.$ (10b)
-Now let us be more specific and consider $F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right)$, where $n>0$ and $a\geq 0$ are some constants, so Eqs. (5) and (9) respectively take the form
+Now let us be more specific and consider
+$F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (11)
+where $n>0$ and $a\geq 0$ are some constants, so Eqs. (5) and (9) respectively take the form
 $u_{\xi\xi}+cu_{\xi}$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (12a)
@@
 One can easily check up that there exists the solution of (12b) satisfying the boundary conditions (10)
-$p=Pu\left(u^{n}-1\right),$ (12)
+$p(u)=p_{0}(u)\equiv Pu\left(u^{n}-1\right),$ (13)
 where
@@
 provided that the speed of the wave is
-$c=P+\frac{a}{P}.$ (14)
+$c=c_{0}\equiv P+\frac{a}{P}.$ (15)
 Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (12a) as
 $u(\xi)=\frac{1}{\left[\exp(nP\xi)+1\right]^{1/n}}.$ (16)
-For the Fisher equation we obtain the well known result ablo
+(An alternative way to obtain (16) was presented in our previous publication kogan4 .)
+For the Fisher equation from (16) follows the well known result ablo
 $u(\xi)=\frac{1}{\left[\exp\left(\xi/\sqrt{6}\right)+1\right]^{2}},$ (17)
@@
 In Appendix A we’ll see how the results (13) - (16) can be generalized in the case when advection is taken into account in addition to diffusion.
+In Appendix C we’ll analyze the solution of (12b) (for $n=1$) for $c$ close to $c_{0}$.
 IV The correct boundary conditions for Eq. (9)
@@
 Note that if for physical reasons $u$ should be strictly nonnegative, the focus solutions (and hence the value of speed $c$ for which $c^{2}\leq 4F(0)$) are unacceptable.
+(Apart from recovering this well known property, we in this paper will say nothing about the front propagation speed selection, an issue which is both important and complicated (saarloos ; ma .)
 Now we understand that the second boundary condition (10b) can be discarded.
@@
 And returning to the previous Section: we did finetune the speed $c$ for the solution of (12b) to be expressed in terms of elementary functions, and not for the solution to exist.
+Acknowledgements.
+We are grateful to G. G. Contreras, A. Gupta, M. Ma, B. Malomed and C. Ou for the insightful discussions.
 Appendix A Generalized Burgers–Huxley equation
@@
 where $\lambda$ is a constant.
-(The Burgers–Huxley equation is known in various fields of applied mathematics.
-For example, it describes transport processes in systems when diffusion and convection are equally important and nonlinear reaction–diffusion processes too.)
+The Burgers–Huxley equation is known in various fields of applied mathematics.
+For example, it describes transport processes in systems when diffusion and convection are equally important and nonlinear reaction–diffusion processes too.
+(An interesting further generalization of the model see in gupta .)
 For the travelling wave solution of (35) we obtain instead of (12b) the equation
@@
 $u_{t}-\lambda u^{n}u_{x}-u_{xx}=u^{n+1}\left(1-u^{n}\right).$ (38)
-Appendix B Perturbation series
+Appendix B Perturbation theory I
 Let us return to Eq.
@@
 $p^{(1)}$ $=F(u),$ (40a)
-$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u}.$ (39b)
+$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u},\hskip 14.22636ptk=2,3,\dots.$ (40b)
 Let us restrict ourselves by the next to leading order approximation with respect to the parameter $1/c$. Substituting thus truncated $p(u)$ into (8) we obtain
@@
 $+\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$ (48)
+Appendix C Perturbation theory II
+In Section III we have solved Eq.
+(12b) for $c=c_{0}$.
+Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series.
+To make life simpler $n=1$, we’ll consider only the case $n=1$.
+Presenting the solution as
+$p(u)=p_{0}(u)+\frac{\delta c}{\sqrt{2}}\delta p(u),$ (49)
+for $\delta p(u)$ in the lowest with respect to $\delta c$ order we obtain equation
+$u(u-1)\delta p_{u}+\left(2u-1+c_{0}\right)\delta p+u(u-1)=0.$ (50)
+Integrating (50) we obtain
+[…]
+where
+[…]
+and $C$ is an arbitrary constant. The lower limits of the integrals in (51) and (52) can be taken arbitrary. Calculating the integral in (52) we get
+$\delta p(u)=\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\left(C-\int\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}\right).$ (53)
+It will be convenient to choose the lower limit in the in integral in (53) to be equal to zero. After that, for $\delta p$ to remain finite when $u\to 0,1$ we should choose $C=0$, thus obtaining
+$\delta p(u)=-\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\int_{0}^{u}\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}.$ (54)
+A pleasant surprise is that $\delta p(u)$, as given by (54), goes to zero when $u\to 0,1$.
+(We couldn’t be sure in that just looking at Eq.
+(50), because the coefficient before $\delta p_{u}$ in the equation goes to zero in the above mentioned limits, hence one can doubt the validity of the linear approximation with respect to $\delta c$ in the vicinity of $u=0$ and $u=1$.
+)

### v4

@@
 $F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (11)
-where $n>0$ and $a\geq 0$ are some constants, so Eqs. (5) and (9) respectively take the form
-$u_{\xi\xi}+cu_{\xi}$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (12a)
-$pp_{u}+cp$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (12b)
-Note that for $n=1/2$ and $a=1$ we recover Fisher equation.
-Equation (12a) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
-One can easily check up that there exists the solution of (12b) satisfying the boundary conditions (10)
+where $n>0$ and $a\geq 0$ are some constants, so (5) takes the form
+$u_{\xi\xi}+cu_{\xi}=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (12)
+Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (12) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
+Starting from (12), we obtain Eq. (9) in the form
+$pp_{u}+cp=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (13)
+One can easily check up that there exists the solution of (13) satisfying the boundary conditions (10)
 $p(u)=p_{0}(u)\equiv Pu\left(u^{n}-1\right),$ (14)
@@
 $c=c_{0}\equiv P+\frac{a}{P}.$ (16)
-Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (12a) as
+Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (12) as
 $u(\xi)=\frac{1}{\left[\exp(nP\xi)+1\right]^{1/n}}.$ (17)
@@
 In Appendix A we’ll see how the results (14) - (17) can be generalized in the case when advection is taken into account in addition to diffusion.
-In Appendix C we’ll analyze the solution of (12b) (for $n=1$) for $c$ close to $c_{0}$.
+In Appendix C we’ll analyze the solution of (13) (for $n=1$) for $c$ close to $c_{0}$.
 IV The correct boundary conditions for Eq. (9)
@@
 of the equation.
-So one may think that we should finetune the speed $c$ for the solution of (12b) to exist, and the solution will not exist for a general value of $c$.
+So one may think that we should finetune the speed $c$ for the solution of (13) to exist, and the solution will not exist for a general value of $c$.
 In fact, didn’t we finetune $c$ in the previous Section?
-However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (12b) for arbitrary $c\gg 1$ is presented.
+However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (13) for arbitrary $c\gg 1$ is presented.
 It looks like we have a problem.
-We have another problem with the boundary conditions (10) - they are exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
-In this Section we’ll solve both problems.
-We’ll show that one of the boundary conditions (10) should be specified, and the other can be discarded.
-Consider the solution of Eq. (5) in the vicinity of $u=0$ and $u=1$. In the first case, the equation can be presented as
-$u_{\xi}$ $=p,$ (18a)
-$p_{\xi}$ $=-cp-F^{\prime}(0)u.$ (18b)
-In the second case, the equation can be presented as
-$v_{\xi}$ $=-p,$ (19a)
-$p_{\xi}$ $=cp-|F^{\prime}(1)|v,$ (19b)
+To solve this problem we have to return to Eq.
+(5) with the boundary conditions (6).
+Actually, the mechanical analogy allows us easily visualize the solution of the equation, but let us forget for the time being about this analogy and look at the equation purely mathematically.
+We immediately realize that the boundary conditions (6) - they are exceptional goursat $u_{\xi\xi}$ is not defined at $\xi=\pm\infty$.
+To make use of the boundary conditions consider Eq.
+(5) for $\xi\to\pm\infty$.
+For $\xi\to+\infty$ the equation can be presented as
+$u_{\xi\xi}+cu_{\xi\xi}+F^{\prime}(0)u=0.$ (19)
+For $\xi\to-\infty$ the equation can be presented as
+$v_{\xi\xi}+cv_{\xi}-|F^{\prime}(1)|v=0,$ (20)
 where $v=1-u$.
-Let us start from analysing (19). The system can be reduced to the canonical form
-$X_{\xi}$ $=s_{1}X,$ (20a)
-$Y_{\xi}$ $=s_{2}Y,$ (20b)
-where
-$X$ $=p+s_{2}v,$ (21a)
-$Y$ $=p+s_{1}v,$ (21b)
-and $s_{1,2}$ are the roots of the characteristic equation
+The solution of (20) is
+$v=Ae^{s_{1}\xi}+Be^{s_{2}\xi},$ (21)
+where $A$ and $B$ are arbitrary constants, and $s_{1,2}$ are the roots of the characteristic equation
 $s^{2}-cs-|F^{\prime}(1)|=0,$ (22)
 that is
-$s_{1,2}$ $=\frac{c\pm\sqrt{c^{2}+4|F^{\prime}(1)|}}{2}.$ (23)
-(Further on $s_{1}$ will be the positive root, and $s_{2}$ the negative.)
-Because $u(\xi)$ should satisfy the boundary condition (6a), we obtain
-$Y=0.$ (24)
-Returning to Eq. (9) we realize that the boundary condition (10a) should be specified as follows: In the vicinity of $u=1$
+$s_{1}$ $=\frac{c+\sqrt{c^{2}+4|F^{\prime}(1)|}}{2},$ (23a)
+$s_{2}$ $=\frac{c-\sqrt{c^{2}+4|F^{\prime}(1)|}}{2}.$ (23b)
+Because of the boundary condition (6a), $B=0$. So this boundary condition should be specified to: For $\xi\to-\infty$
+$1-u\sim e^{s_{1}\xi}.$ (24)
+Now let us come to (19). The characteristic equation is
+$s^{2}+cs+F^{\prime}(0)=0,$ (25)
+that is
+$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4F^{\prime}(0)}}{2}.$ (26)
+We have either the real roots (both being negative), the double (negative) root, or the complex roots (both with the negative real part). In the first case solution of (19) is
+$u=Ae^{s_{3}\xi}+Be^{s_{4}\xi},$ (27)
+for the case of double root the solution is
+$u=e^{-c\xi/2}(A+B\xi),$ (28)
+and in the complex case the solution is
+$u=e^{-c\xi/2}\left[A\cos(\alpha\xi)+B\sin(\alpha\xi)\right],$ (29)
+where $\alpha=[4F^{\prime}(0)-c^{2}]/2$. In any case the solution of (19) satisfies the boundary condition (6b) for any $A$ and $B$. So the boundary condition (6b) can be discarded.
+Because we have lost one boundary condition the solution of (5) contains an arbitrary constant.
+And here our mechanical analogy helps us to understand that this (irrelevant) constant describes the translation of the solution in ”time” $\xi$.
+After we understood the true nature of the boundary conditions (6) for (5), the situation with Eq.
+(9) becomes obvious.
+Note that the boundary conditions (10) are also exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
+However, looking back at (24) we realize that the boundary condition (10a) should be specified as follows: For $u\to 1$
 $p=s_{1}(u-1).$ (30)
@@
 However, if we want, for example, to solve (9) numerically, the condition (30) should be used from the beginning.
-Now let us come to (18). The characteristic equation for the system is
-$s^{2}+cs+F^{\prime}(0)=0.$ (26)
-The roots of the characteristic equations are
-$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4F^{\prime}(0)}}{2}.$ (27)
-We have to consider separately the case of the real roots, the case of the complex roots and the case of the double root.
-Let us start from the case of the real roots ($c^{2}>4F^{\prime}(0)$). In this case we can reduce the system (18) to
-$\frac{dX}{s_{3}X}=\frac{dY}{s_{4}Y},$ (28)
-where
-$X$ $=p+s_{4}v,$ (29a)
-$Y$ $=p+s_{3}v.$ (29b)
-The general integral is given by the equation
-$Y=CX^{\frac{s_{4}}{s_{3}}}.$ (30)
-Because $s_{3}$ and $s_{4}$ have the same sign, $X$ approaches zero with $Y$, and all the integral curves pass through the origin, which is a node.
-(Note that $c$ given by (15) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
-For $c^{2}<4F^{\prime}(0)$ the characteristic equation has two conjugate imaginary roots $\alpha+\beta i,\alpha-\beta i$, $\beta\neq 0$) we can reduce the system (18) to the form goursat
-$\frac{dX}{\alpha X-\beta Y}=\frac{dY}{\beta X+\alpha Y}$ (31)
-from which we derive
-$\frac{XdX+YdY}{\alpha\left(X^{2}+Y^{2}\right)}=\frac{XdY-YdX}{\beta\left(X^{2}+Y^{2}\right)}$ (32)
-The general integral of the equation (18) is therefore represented by the equation
-$\sqrt{X^{2}+Y^{2}}=Ce^{\frac{\alpha}{\beta}\tan^{-1}\frac{Y}{X}}.$ (33)
-All these curves have the form of spirals which approach the origin as an asymptotic point. The origin is said to he a focus.
-If the characteristic equation has a double root $s=c/2$, which occurs for $c^{2}=4F^{\prime}(0)$, the system (18) reduces to the form
-$\frac{dX}{X}=\frac{dY}{X+Y},$ (34)
-and the general integral is $Y=CX+X\ln X$.
-In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e^{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$.
-When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus.
-Note that if for physical reasons $u$ should be strictly nonnegative, the focus solutions (and hence the value of speed $c$ for which $c^{2}\leq 4F(0)$) are unacceptable.
-(Apart from recovering this well known property, we in this paper will say nothing about the front propagation speed selection, an issue which is both important and complicated (saarloos ; ma .)
-Now we understand that the second boundary condition (10b) can be discarded.
-It is satisfied automatically.
-And returning to the previous Section: we did finetune the speed $c$ for the solution of (12b) to be expressed in terms of elementary functions, and not for the solution to exist.
+And looking back at our analysis of (19) we realize that any solution of Eq. (9) will satisfy (10b), so the second boundary condition (10b) can be discarded. It is satisfied automatically.
+So finally, in the previous Section we were solving the first order ODE with the single boundary condition.
+The solution should exist for any given value of the speed $c$ and shouldn’t contain an undefined constant.
+We did finetune the speed $c$ in the previous Section for the solution of (13) to be expressed in terms of elementary functions, and not for the solution to exist.
+Our analysis of (19) allows us to understand (partially) the analytic properties of $p(u)$.
+If $c^{2}\geq 4F^{\prime}(0)$, from (27) and (28) we can expect that $p(u)$ would be either a single - or a double-valued function.
+(Further analysis is necessary to decide which option is realised.) Note that $c$ given by (16) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
+Looking at (29) (which corresponds to $c^{2}<4F^{\prime}(0)$) on the other hand, we realize that $p(u)$ would be an infinitely-valued function.
+Note that if for physical reasons $u$ should be strictly nonnegative, such values of speed are unacceptable, because from (29) follows that $u$ would acquire negative values in the vicinity of $\xi=+\infty$.
+(Apart from recovering this well known property, we in this paper will say nothing about the front propagation speed selection, an issue which is both important and complicated saarloos ; ma .)
 Acknowledgements.
-We are grateful to G. G. Contreras, A. Gupta, M. Ma, B. Malomed and C. Ou for the insightful discussions.
+We are grateful to G. G. Contreras, A. Gupta, M. Ma, B. Malomed, C. Ou and R. Selvaraj for the insightful discussions.
 Appendix A Generalized Burgers–Huxley equation
-The results obtained in Section II for the generalized KPP equation (12a) can be easily generalised to the generalized Burgers–Huxley equation cherniha ; kushner
+The results obtained in Section II for the generalized KPP equation (12) can be easily generalised to the generalized Burgers–Huxley equation cherniha ; kushner
 $u_{t}-\lambda u^{n}u_{x}-u_{xx}=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (31)
@@
 (An interesting further generalization of the model see in gupta .)
-For the travelling wave solution of (35) we obtain instead of (12b) the equation
+For the travelling wave solution of (31) we obtain instead of (13) the equation
 $pp_{u}+\lambda u^{n}p+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (32)
@@
 In Section III we have solved Eq.
-(12b) for $c=c_{0}$.
+(13) for $c=c_{0}$.
 Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series.

### v5

@@
 We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed.
-The solutions are written in terms of elementary functions.
+Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
 I Introduction
@@
 where $F$ is a sufficiently smooth function with the properties that $F(0)=F(1)=0$, $F^{\prime}(0)>0$ and $F^{\prime}(1)<0$.
-II Travelling waves
 Equation (3) admits traveling wave solutions
@@
 The motion is with friction, and the coefficient of friction is equal to $c$. The point $u=0$ is the point of stable equilibrium, and the point $u=1$ unstable.
-III The ODE which doesn’t contain explicitly the independent variable
-Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea benguria ; kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
+Equation (5) is a nonlinear nonintegrable differential equation.
+The Cauchy problem cannot be solved by the inverse scattering transform for this equation and the problem of finding some exact solutions for this equation is an important task.
+Several methods of obtaining such solutions has been developped.
+We will mention here a few of them: the tanh-function method 4 ; 5 ; 6 , the Jacobi elliptic function expansion method 8 , the simplest equation method 11 ; 12 , the Exp-function method 14 , the $G^{\prime}/G$ expansion method 15 ; 16 ; 17 ; 18 , and the Kudryashov method 19 .
+In this paper we will use the method of exclusion of the independent variable benguria ; kogan3 ; kogan4 ; kogan5 , which allows to reduce the problem of integration of the second order ODE to the problem of integration of the first order ODE, to obtain the exact solutions of Eq.
+(5).
+Thus we were able to obtain the exact solutions of the KPP equation which to the best of our knowledge were not known before.
+Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
+II The ODE which doesn’t contain explicitly the independent variable
+Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea to consider $u$ as the new independent variable and
 $p=\frac{du}{d\xi}$ (8)
@@
 In Appendix C we’ll analyze the solution of (13) (for $n=1$) for $c$ close to $c_{0}$.
-IV The correct boundary conditions for Eq. (9)
+III The correct boundary conditions for Eq. (9)
 In the previous Section we postulated two boundary conditions (10) for (9).
@@
 And here our mechanical analogy helps us to understand that this (irrelevant) constant describes the translation of the solution in ”time” $\xi$.
-After we understood the true nature of the boundary conditions (6) for (5), the situation with Eq.
+After we understood the true nature of the boundary conditions (6) for (5), the situation with the boundary conditions (10) for Eq.
 (9) becomes obvious.
 Note that the boundary conditions (10) are also exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
-However, looking back at (24) we realize that the boundary condition (10a) should be specified as follows: For $u\to 1$
+In the vicinity of $u=1$, Eq. (9) can be simplified to
+$qq_{v}-cq=|F^{\prime}(1)|v,$ (30)
+where again $v=1-u$, and $q\equiv dv/d\xi$. Equation (30) has two solutions
+$q$ $=s_{1}v,$ (31a)
+$q$ $=s_{2}v,$ (31b)
+where $s_{1,2}$ are given by (23). Only the solution (31a) satisfies the condition (6a). Hence we realize that the boundary condition (10a) should be specified as follows: For $u\to 1$
 $p=s_{1}(u-1).$ (32)
-For practical purposes we can sometimes continue to use this boundary condition in the previous form (10a), like we deed in Section III.
-The differential equation itself produces the correct condition (30).
-However, if we want, for example, to solve (9) numerically, the condition (30) should be used from the beginning.
-And looking back at our analysis of (19) we realize that any solution of Eq. (9) will satisfy (10b), so the second boundary condition (10b) can be discarded. It is satisfied automatically.
+(Not the difference between the proportionality sign in (24) and the equality sign in (32).) One can easily check up that our solution (14) satisfies this boundary condition.
+In the vicinity of $u=0$, Eq. (9) can be simplified to
+$pp_{v}+cp=F^{\prime}(0)u.$ (33)
+Equation (33) has two solutions
+$p$ $=s_{3}u,$ (34a)
+$p$ $=s_{4}u,$ (34b)
+where $s_{3,4}$ are given by (26).
+Because both $s_{3}$ and $s_{4}$ have negative real part, both solutions will satisfy the boundary condition (6b).
+So we understood that the second boundary condition (10b) can be discarded – it is satisfied automatically.
 So finally, in the previous Section we were solving the first order ODE with the single boundary condition.
@@
 Appendix C Perturbation theory II
-In Section III we have solved Eq.
+In Section II we have solved Eq.
 (13) for $c=c_{0}$.

### v6

@@
 We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed.
-Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
+Additionally we obtain the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
 I Introduction
@@
 Equation (1) arises in many physical, biological, and chemical problems involving diffusion and nonlinear growth.
-It is convenient to introduce the nondimensional quantities $x^{*},t^{*},u^{*}$, so after dropping the asterisks, equation (1) takes the nondimensional form
+It is convenient to introduce the nondimensional quantities
+$x^{*}$ $=x\sqrt{k/\nu},$ (2a)
+$t^{*}$ $=kt,$ (2b)
+$u^{*}$ $=u/\kappa,$ (2c)
+so after dropping the asterisks, equation (1) takes the nondimensional form
 $u_{t}-u_{xx}=u\left(1-u\right).$ (3)
@@
 $u(x,t)=u(\xi),\hskip 14.22636pt\xi=x-ct,$ (5)
-where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
-$u_{\xi\xi}+cu_{\xi}=F(u)$ (5)
+where $c$ is the wave speed (measured in the units of $\sqrt{\nu/k}$), and $u(\xi)$ satisfies the equation
+$u_{\xi\xi}+cu_{\xi}=-F(u)$ (6)
 with the boundary conditions
@@
 It describes motion of the Newtonian particle with the mass 1 ($u$ being the coordinate of the particle and $\xi$ the time) in the potential well $U(u)$ defined by the equation
-$\frac{dU}{du}=-F(u).$ (7)
+$\frac{dU}{du}=F(u).$ (8)
 The motion is with friction, and the coefficient of friction is equal to $c$. The point $u=0$ is the point of stable equilibrium, and the point $u=1$ unstable.
@@
 The Cauchy problem cannot be solved by the inverse scattering transform for this equation and the problem of finding some exact solutions for this equation is an important task.
-Several methods of obtaining such solutions has been developped.
+Several methods of obtaining such solutions has been developed.
 We will mention here a few of them: the tanh-function method 4 ; 5 ; 6 , the Jacobi elliptic function expansion method 8 , the simplest equation method 11 ; 12 , the Exp-function method 14 , the $G^{\prime}/G$ expansion method 15 ; 16 ; 17 ; 18 , and the Kudryashov method 19 .
-In this paper we will use the method of exclusion of the independent variable benguria ; kogan3 ; kogan4 ; kogan5 , which allows to reduce the problem of integration of the second order ODE to the problem of integration of the first order ODE, to obtain the exact solutions of Eq.
+In this paper we will use the method of exclusion of the independent variable benguria ; rosu ; kogan3 ; kogan4 ; kogan5 , which allows to reduce the problem of integration of the second order ODE to the problem of integration of the first order ODE, to obtain the exact solutions of Eq.
 (6).
 Thus we were able to obtain the exact solutions of the KPP equation which to the best of our knowledge were not known before.
-Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
+Additionally we obtain the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
 II The ODE which doesn’t contain explicitly the independent variable
 Equation (6) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea to consider $u$ as the new independent variable and
-$p=\frac{du}{d\xi}$ (8)
+$p=u_{\xi}$ (9)
 as the new dependent variable. In the new variables (6) takes the form of Abel equation of the second kind polyanin .
-$pp_{u}+cp=F(u).$ (9)
+$pp_{u}+cp=-F(u).$ (10)
 The boundary conditions for (6) were (7). The boundary conditions for (10) are (however, see the next Section)
@@
 $F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (12)
-where $n>0$ and $a\geq 0$ are some constants, so (5) takes the form
-$u_{\xi\xi}+cu_{\xi}=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (12)
-Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (12) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
-Starting from (12), we obtain Eq. (9) in the form
-$pp_{u}+cp=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (13)
-One can easily check up that there exists the solution of (13) satisfying the boundary conditions (10)
+where $n>0$ and $a\geq 0$ are some constants, so (6) will take the form
+$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (13)
+Note that for $n=1/2$ and $a=1$ we recover Fisher equation.
+Equation (13) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
+Starting from (13), we obtain Eq.
+(10) in the form
+$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (14)
+In search of elementary solutions of (14) let us return to Eq.
+(10), and in accordance with the well-known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us even generalize it to
+$pp_{u}+\gamma(u)p=-F(u),$ (15)
+where $\gamma(u)$ is some function. Being inspired by the method of factorization gonzales , we present the r.h.s. of (15) as
+$-F(u)=\Phi(u)\Psi(u).$ (16)
+After that we realize that there exists the particular solution of (15)
+$p(u)=P\Phi(u),$ (17)
+provided the unknown multiplier $P$ and the function $\gamma(u)$ satisfy equation
+$P^{2}\Phi_{u}+P\gamma(p)=\Psi(u).$ (18)
+This includes the case of nonlinear $F(u)$ considered in Ref. an
+$F(U)=F^{\prime}(0)\left[u-A(u)\right]\left[1+\chi A^{\prime}(u)\right],$ (19)
+where $A(u)$ is some function, $\chi$ is some numerical parameter, and $\gamma(p)=c=$ const. In fact, taking
+$\Phi(u)=F^{\prime}(0)\left[A(u)-u\right]$ (20)
+we obtain (18) in the form
+$P^{2}F^{\prime}(0)\left[A^{\prime}(u)-1\right]+Pc=1+\chi A^{\prime}(u).$ (21)
+We see that choosing $P$ and $c$ we can turn (21) into the identity.
+Now let us return to Eq. (14) and take
+$\Phi(u)=u\left(u^{n}-1\right),\hskip 28.45274pt\Psi(u)=\left(u^{n}+a\right).$ (22)
+Equation (18) takes the form
+$P^{2}\left[(n+1)u^{n}-1\right]+Pc=u^{n}+a.$ (23)
+We immediately realize that the solution of (14) satisfying the boundary conditions (11)
 $p(u)=p_{0}(u)\equiv Pu\left(u^{n}-1\right),$ (24)
@@
 $P=\frac{1}{\sqrt{n+1}},$ (25)
-provided that the speed of the wave is
+exists if the speed of the wave is
 $c=c_{0}\equiv P+\frac{a}{P}.$ (26)
-Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (12) as
+Substituting thus obtained $p(u)$ into (9) and integrating we obtain the solution of (13)
 $u(\xi)=\frac{1}{\left[\exp(nP\xi)+1\right]^{1/n}}.$ (27)
-(An alternative way to obtain (17) was presented in our previous publication kogan4 .)
 For the Fisher equation from (27) follows the well known result ablo
@@
 and the condition (26) takes the form $c=5/\sqrt{6}$.
-In Appendix A we’ll see how the results (14) - (17) can be generalized in the case when advection is taken into account in addition to diffusion.
-In Appendix C we’ll analyze the solution of (13) (for $n=1$) for $c$ close to $c_{0}$.
+To additionally illustrate the method of integration used in this Section, in the Appendix A we obtain in the framework of the method the exact solution of the generalized Duffing-Van der Pol equation.
+In Appendix B we’ll see how the results (24) - (27) can be generalized in the case when advection is taken into account in addition to diffusion.
+The connection between the existence of the elementary solution of (10) and the possible symmetry of the equation will be analysed in our next publication.
 III The correct boundary conditions for Eq. (10)
@@
 However, the equation is of the first order.
-The first impression is that the problem is overdetermined, and the solution obtained there exists due to the finetuning of the l.h.s.
+The first impression is that the problem is overdetermined, and the existence of the solution of (10) demands the finetuning of the l.h.s.
 of the equation.
-So one may think that we should finetune the speed $c$ for the solution of (13) to exist, and the solution will not exist for a general value of $c$.
-In fact, didn’t we finetune $c$ in the previous Section?
-However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (13) for arbitrary $c\gg 1$ is presented.
+Being more specific, one may think that the solution of (14) will not exist for a general value of $c$.
+However, the last statement immediately comes into contradiction with the results of the Appendix C, where the solution of (14) for arbitrary $c\gg 1$ is presented.
 It looks like we have a problem.
@@
 (6) with the boundary conditions (7).
-Actually, the mechanical analogy allows us easily visualize the solution of the equation, but let us forget for the time being about this analogy and look at the equation purely mathematically.
-We immediately realize that the boundary conditions (6) - they are exceptional goursat $u_{\xi\xi}$ is not defined at $\xi=\pm\infty$.
-To make use of the boundary conditions consider Eq.
-(5) for $\xi\to\pm\infty$.
-For $\xi\to+\infty$ the equation can be presented as
-$u_{\xi\xi}+cu_{\xi\xi}+F^{\prime}(0)u=0.$ (19)
-For $\xi\to-\infty$ the equation can be presented as
+The boundary conditions are exceptional goursat $u_{\xi\xi}$ is not defined at $\xi=\pm\infty$.
+However, for $\xi\to+\infty$ the equation asymptotically becomes
+$u_{\xi\xi}+cu_{\xi\xi}+F^{\prime}(0)u=0,$ (29)
+and for $\xi\to-\infty$
 $v_{\xi\xi}+cv_{\xi}-|F^{\prime}(1)|v=0,$ (30)
@@
 And here our mechanical analogy helps us to understand that this (irrelevant) constant describes the translation of the solution in ”time” $\xi$.
-After we understood the true nature of the boundary conditions (6) for (5), the situation with the boundary conditions (10) for Eq.
-(9) becomes obvious.
-Note that the boundary conditions (10) are also exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
-In the vicinity of $u=1$, Eq. (9) can be simplified to
+After we understood the true nature of the boundary conditions (7) for (6), the situation with the boundary conditions (11) for Eq. (10) - which are also exceptional - becomes obvious.
+In the vicinity of $u=1$, Eq. (10) becomes asimptotically
 $qq_{v}-cq=|F^{\prime}(1)|v,$ (40)
@@
 (Not the difference between the proportionality sign in (34) and the equality sign in (42).) One can easily check up that our solution (24) satisfies this boundary condition.
-In the vicinity of $u=0$, Eq. (9) can be simplified to
+In the vicinity of $u=0$, Eq. (10) becomes asimptotically
 $pp_{v}+cp=F^{\prime}(0)u.$ (43)
@@
 Acknowledgements.
-We are grateful to G. G. Contreras, A. Gupta, M. Ma, B. Malomed, C. Ou and R. Selvaraj for the insightful discussions.
-Appendix A Generalized Burgers–Huxley equation
+We are grateful to G. Gonzales, A. Gupta, M. Ma, B. Malomed, A. Mogilner, H. C. Rosu, C. Ou and R. Selvaraj for the insightful discussions.
+Appendix A Exact solution of the generalized Duffing –Van der Pol equation
+Consider the generalized Duffing –Van der Pol equation
+$X_{tt}+\left(a-bF_{X}\right)X_{t}=F(X),$ (45)
+where $X(t)$ is a deviation from the balance of the oscillator, $t$ is time, $F(X)$ is the force field determined by the model, and $a,b$ are the parameters determining the generalized friction.
+(Equation (45) describes, in particular, travelling waves in a lossy transmission line with nonlinear capacitance kogan3 .) Introducing new dependent variable
+$p=X_{t}$ (46)
+and considering $X$ as the independent variable, we reduce (45) to the first order ODE
+$pp_{X}+\left(a-bF_{X}\right)p=F(X).$ (47)
+The particular solution of (47)
+$p(X)=bF(X)$ (48)
+exists if
+$ab=1.$ (49)
+Looking back at Section II we realize that the solution corresponds to the factorization
+$F(X)=F(X)\cdot 1.$ (50)
+Substituting thus obtained function $p(X)$ into (46) we obtain the solution of (45)
+$t=\int\frac{dX}{bF(X)}.$ (51)
+Consider $F(X)$ which has two adjacent zeros at $X_{1}$ and $X_{2}$ ($F(X_{1})=F(X_{2})=0$), one of these points being the point of maximum of the potential $V(X)$ defined by the force field $F(X)$, the other - the point of minimum.
+Then (51) describes explicitly the motion of the particle in the potential well $V(X)$ which starts in the infinite past at the potential maximum and ends in the infinite future at the potential minimum.
+Finetuning of the friction (Eq.
+(49)) is necessary for the absence of oscillations in the vicinity of the potential minimum, which gives the opportunity to integrate (45) in quadratures.
+Let us be even more specific and consider the Duffing –Van der Pol equation from Ref. 19
+$X_{tt}+\left(\alpha+\beta X^{2}\right)X_{t}=gX-X^{3}.$ (52)
+Equation (52) is the particular case of (45) corresponding to
+$F(X)$ $=gX-X^{3},$ (53a)
+$a$ $=\alpha+\beta g/3$ (53b)
+$b$ $=\beta/3.$ (53c)
+Thus the condition (49) takes the form
+$\alpha=3/\beta-\beta g/3;$ (54)
+integration in (51) can be easily performed and we reproduce the result of Ref. 19
+$X(t)=\frac{\sqrt{g}}{\sqrt{1-e^{-2\beta gt/3}}}.$ (55)
+Appendix B Generalized Burgers–Huxley equation
 The results obtained in Section II for the generalized KPP equation (13) can be easily generalised to the generalized Burgers–Huxley equation cherniha ; kushner
@@
 $u_{t}-\lambda u^{n}u_{x}-u_{xx}=u^{n+1}\left(1-u^{n}\right).$ (59)
-Appendix B Perturbation theory I
+Appendix C Perturbation theory I
 Let us return to Eq.
@@
 $+\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$ (69)
-Appendix C Perturbation theory II
+Appendix D Perturbation theory II
 In Section II we have solved Eq.
 (13) for $c=c_{0}$.
-Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series.
+Motivated by the perturbation theory of the previous Section, let us study the solution of (13) for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series.
 To make life simpler $n=1$, we’ll consider only the case $n=1$.
-Presenting the solution as
+Presenting the solution of (14) (for $n=1$) as
 $p(u)=p_{0}(u)+\frac{\delta c}{\sqrt{2}}\delta p(u),$ (70)

### v7

@@
 $p(0)$ $=0.$ (11b)
-Now let us be more specific and consider
-$F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (12)
-where $n>0$ and $a\geq 0$ are some constants, so (6) will take the form
-$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (13)
-Note that for $n=1/2$ and $a=1$ we recover Fisher equation.
-Equation (13) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
-Starting from (13), we obtain Eq.
-(10) in the form
-$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (14)
-In search of elementary solutions of (14) let us return to Eq.
-(10), and in accordance with the well-known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us even generalize it to
+In search of elementary solutions of (10), in accordance with the well-known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us generalize it to
 $pp_{u}+\gamma(u)p=-F(u),$ (12)
@@
 This includes the case of nonlinear $F(u)$ considered in Ref. an
-$F(U)=F^{\prime}(0)\left[u-A(u)\right]\left[1+\chi A^{\prime}(u)\right],$ (19)
-where $A(u)$ is some function, $\chi$ is some numerical parameter, and $\gamma(p)=c=$ const. In fact, taking
-$\Phi(u)=F^{\prime}(0)\left[A(u)-u\right]$ (20)
+$F(U)=\left[u-A(u)\right]\left[\chi A^{\prime}(u)+a\right],$ (16)
+where $A(u)$ is some function, $\chi,a$ are some numerical parameters, and $\gamma(p)=c=$ const. In fact, taking
+$\Phi(u)=\left[A(u)-u\right]$ (17)
 we obtain (15) in the form
-$P^{2}F^{\prime}(0)\left[A^{\prime}(u)-1\right]+Pc=1+\chi A^{\prime}(u).$ (21)
-We see that choosing $P$ and $c$ we can turn (21) into the identity.
-Now let us return to Eq. (14) and take
-$\Phi(u)=u\left(u^{n}-1\right),\hskip 28.45274pt\Psi(u)=\left(u^{n}+a\right).$ (22)
-Equation (18) takes the form
+$P^{2}\left[A^{\prime}(u)-1\right]+Pc=\chi A^{\prime}(u)+a.$ (18)
+Choosing $P^{2}=\chi$ and $c=P+a/P$ we turn (18) into identity.
+Now let us be more specific and consider
+$F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (19)
+where $n>0$ and $a\geq 0$ are some constants, so that (6) will take the form
+$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (20)
+Note that for $n=1/2$ and $a=1$ we recover Fisher equation.
+Equation (20) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
+Starting from (20), we obtain Eq.
+(10) in the form
+$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (21)
+We can take
+$\Phi(u)=u\left(u^{n}-1\right),\hskip 28.45274pt\Psi(u)=\left(u^{n}+a\right),$ (22)
+after which (15) takes the form
 $P^{2}\left[(n+1)u^{n}-1\right]+Pc=u^{n}+a.$ (23)
@@
 After we understood the true nature of the boundary conditions (7) for (6), the situation with the boundary conditions (11) for Eq. (10) - which are also exceptional - becomes obvious.
-In the vicinity of $u=1$, Eq. (10) becomes asimptotically
+In the vicinity of $u=1$, Eq. (10) becomes asymptotically
 $qq_{v}-cq=|F^{\prime}(1)|v,$ (40)
@@
 (Not the difference between the proportionality sign in (34) and the equality sign in (42).) One can easily check up that our solution (24) satisfies this boundary condition.
-In the vicinity of $u=0$, Eq. (10) becomes asimptotically
-$pp_{v}+cp=F^{\prime}(0)u.$ (43)
-Equation (43) has two solutions
+In the vicinity of $u=0$, Eq. (10) becomes asymptotically
+$pp_{u}+cp=-F^{\prime}(0)u.$ (43)
+Equation (43) has two solutions (we’ll consider only the case $c^{2}>4F^{\prime}(0)$)
 $p$ $=s_{3}u,$ (44a)
@@
 where $s_{3,4}$ are given by (36).
-Because both $s_{3}$ and $s_{4}$ have negative real part, both solutions will satisfy the boundary condition (7b).
+Because both $s_{3}$ and $s_{4}$ are negative, both solutions will satisfy the boundary condition (7b).
 So we understood that the second boundary condition (11b) can be discarded – it is satisfied automatically.
+Note that for the case considered explicitly in the previous Section ($F(u)$ being given by (19) and $c$ being given by (26), obviously fulfilling the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$) we get
+$s_{3}$ $=-\frac{1}{\sqrt{n+1}},$ (45a)
+$s_{4}$ $=-a\sqrt{n+1}.$ (45b)
+Looking at Eq. (24) we realise that the solution (44a) is relevant in this case.
 So finally, in the previous Section we were solving the first order ODE with the single boundary condition.
@@
 If $c^{2}\geq 4F^{\prime}(0)$, from (37) and (38) we can expect that $p(u)$ would be either a single - or a double-valued function.
-(Further analysis is necessary to decide which option is realised.) Note that $c$ given by (26) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
+(Further analysis is necessary to decide which option is realised.)
 Looking at (39) (which corresponds to $c^{2}<4F^{\prime}(0)$) on the other hand, we realize that $p(u)$ would be an infinitely-valued function.
@@
 where
-$p^{(1)}$ $=F(u),$ (61a)
+$p^{(1)}$ $=-F(u),$ (62a)
 $p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u},\hskip 14.22636ptk=2,3,\dots.$ (62b)
 Let us restrict ourselves by the next to leading order approximation with respect to the parameter $1/c$. Substituting thus truncated $p(u)$ into (9) we obtain
-$\frac{du}{d\xi}=\frac{1}{c}F(u)\left(1-\frac{1}{c}\frac{dF(u)}{du}\right)$ (62)
+$\frac{du}{d\xi}=-\frac{1}{c}F(u)\left(1+\frac{1}{c}\frac{dF(u)}{du}\right)$ (63)
 or, equivalently (with the accepted precision),
-$\frac{d\xi}{c}=\frac{du}{F(u)}+\frac{1}{c}\frac{dF(u)}{F(u)}.$ (63)
+$\frac{d\xi}{c}=-\frac{du}{F(u)}+\frac{1}{c}\frac{dF(u)}{F(u)}.$ (64)
 Integrating (64) we get
-$\frac{\xi}{c}=\int\frac{du}{F(u)}+\frac{1}{c}\ln|F(u)|.$ (64)
+$\frac{\xi}{c}=-\int\frac{du}{F(u)}+\frac{1}{c}\ln|F(u)|.$ (65)
 To make our life simpler, further on consider only the particular case of $F(u)$ corresponding to the Fisher equation:
@@
 (21) for $c=c_{0}$.
-Motivated by the perturbation theory of the previous Section, let us study the solution of (13) for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series.
-To make life simpler $n=1$, we’ll consider only the case $n=1$.
-Presenting the solution of (14) (for $n=1$) as
-$p(u)=p_{0}(u)+\frac{\delta c}{\sqrt{2}}\delta p(u),$ (70)
-for $\delta p(u)$ in the lowest with respect to $\delta c$ order we obtain equation
-$u(u-1)\delta p_{u}+\left(2u-1+c_{0}\right)\delta p+u(u-1)=0.$ (71)
+Let us now study the solution of the equation for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the linear with respect to $\delta c$ approximation.
+To make life simpler, we’ll consider only the case $n=1$.
+Presenting the solution (for $n=1$) as
+$p(u)=p_{0}(u)+\frac{\delta c}{\sqrt{2}}p_{1}(u)$ (71)
+and substituting (71) into (21), for $p_{1}$ in linear approximation we obtain equation
+$u(u-1){p_{1}}_{u}+\left(2u-1+c_{0}\right)p_{1}+u(u-1)=0.$ (72)
 Integrating (72) we obtain
@@
 and $C$ is an arbitrary constant. The lower limits of the integrals in (73) and (74) can be taken arbitrary. Calculating the integral in (74) we get
-$\delta p(u)=\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\left(C-\int\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}\right).$ (74)
-It will be convenient to choose the lower limit in the in integral in (74) to be equal to zero. After that, for $\delta p$ to remain finite when $u\to 0,1$ we should choose $C=0$, thus obtaining
-$\delta p(u)=-\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\int_{0}^{u}\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}.$ (75)
-A pleasant surprise is that $\delta p(u)$, as given by (75), goes to zero when $u\to 0,1$.
-(We couldn’t be sure in that just looking at Eq.
-(71), because the coefficient before $\delta p_{u}$ in the equation goes to zero in the above mentioned limits, hence one can doubt the validity of the linear approximation with respect to $\delta c$ in the vicinity of $u=0$ and $u=1$.
-)
+$p_{1}(u)=\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\left(C-\int\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}\right).$ (75)
+Taking into account the boundary condition $p_{1}(1)=0$ we obtain
+$p_{1}(u)=\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\int_{u}^{1}\frac{\left(1-v\right)^{c_{0}+1}dv}{v^{c_{0}-1}}.$ (76)
+A question of the validity of the linear approximation presented above is quite delicate, because the coefficient before ${p_{1}}_{u}$ in Eq.
+(72) goes to zero when $u\to 0,1$.
+A good sign is that $p_{1}$ given by (76) has the asymptotic behavior $p_{1}\sim(u-1)$ in the vicinity of $u=1$, in accordance with (42).
+However, in the vicinity of $u=0$, $p_{1}$ has the asymptotic behavior $p_{1}\sim u$, as it should be in accordance with (44), only for $c_{0}>2$.
+(For $c_{0}<2$, $p_{1}\sim u^{c_{0}-1}$.) So we would like to believe that for $c_{0}>2$ the linear approximation is valid everywhere (for $0\leq u\leq 1$), and for $c_{0}<2$ the linear approximation is valid everywhere apart from the vicinity of $u=0$.

### v8

@@
 Additionally we obtain the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
+Keywords: Reaction-diffusion equation; Fisher equation; travelling waves; exact solutions.
 I Introduction
@@
 where $F$ is a sufficiently smooth function with the properties that $F(0)=F(1)=0$, $F^{\prime}(0)>0$ and $F^{\prime}(1)<0$.
+Fisher-KPP equation belongs to the class of reaction-diffusion equations: in fact, it is one of the simplest semilinear reaction-diffusion equations, which can exhibit traveling wave solutions that switch between equilibrium states given by $F(u)=0$.
+Such equations occur, e.g., in ecology, where it describes, like it was mentioned above, the population growth, physiology, combustion, crystallization, plasma physics, and in general phase transition problems.
 Equation (4) admits traveling wave solutions
@@
 Several methods of obtaining such solutions has been developed.
-We will mention here a few of them: the tanh-function method 4 ; 5 ; 6 , the Jacobi elliptic function expansion method 8 , the simplest equation method 11 ; 12 , the Exp-function method 14 , the $G^{\prime}/G$ expansion method 15 ; 16 ; 17 ; 18 , and the Kudryashov method 19 .
+We will mention here a few of them: the tanh-function method 4 ; 5 ; 6 , the Jacobi elliptic function expansion method 8 , the simplest equation method 11 ; 12 , the Exp-function method 14 , the $G^{\prime}/G$ expansion method 15 ; 16 ; 17 ; 18 , the Kudryashov method 19 , and the attached flow method iacobescu ; polyanin .
 In this paper we will use the method of exclusion of the independent variable benguria ; rosu ; kogan3 ; kogan4 ; kogan5 , which allows to reduce the problem of integration of the second order ODE to the problem of integration of the first order ODE, to obtain the exact solutions of Eq.
@@
 (Apart from recovering this well known property, we in this paper will say nothing about the front propagation speed selection, an issue which is both important and complicated saarloos ; ma .)
+IV Conclusions
+We have obtained exact travelling wave solutions of the KPP equation for a wide class of the reaction terms and similar for the generalized Burgers–Huxley equation.
+Our method of integration of the second order differential equations is based on the choice of the original dependent variable as the new independent variable, and the derivative of the original dependent variable - as the new dependent variable.
+The boundary conditions for thus obtained first order differential equations were thoroughly studied.
+All this gave us the opportunity to obtain the exact solutions mentioned above.
 Acknowledgements.